# High-order Computation of Floquet Multipliers and Subspaces using Multistep Methods

### v1

@@ Abstract
Accurate and efficient computation of Floquet multipliers and subspaces is essential for analyzing limit cycle in dynamical systems and periodic steady state in Radio Frequency (RF) simulation. This problem is typically addressed by solving a periodic linear eigenvalue problem, which is discretized from the linear periodic time-varying system using one-step methods. The backward Euler method offers a computationally inexpensive overall workflow but has limited accuracy. In contrast, one-step collocation methods achieve higher accuracy through over-sampling, explicit matrix construction, and condensation, thus become costly for large-scale sparse cases. We apply multistep methods to derive a periodic polynomial eigenvalue problem, which introduces additional spurious eigenvalues. Under mild smoothness assumptions, we prove that as the stepsize decreases, the computed Floquet multipliers and their associated invariant subspace converge with higher order, while the spurious eigenvalues converge to zero. To efficiently solve large-scale problems, we propose pTOAR, a memory-efficient iterative algorithm for computing the dominant Floquet eigenpairs. Numerical experiments demonstrate that multistep methods achieves high order accuracy, while its computational and memory costs are only marginally higher than those of the backward Euler method.
@@ 1 Introduction
A linear periodic time-varying (LPTV) system is described by eq.˜1:
$\frac{\mathrm{d}x(t)}{\mathrm{d}t}=G(t)x(t),$ (1)
@@
where $G(t)\in\mathbb{R}^{n\times n}$ is a $T$-periodic matrix function, i.e., $G(t+T)=G(t)$, and $x(t)\in\mathbb{R}^{n}$ is the state vector. The state transition matrix $\Phi(T,0)$ of eq.˜1 is obtained by integrating eq.˜1 from $[0,T]$ with initial value matrix $X(0)=I_{n}$. The Floquet multipliers and eigenvectors are defined as the eigenpairs of $\Phi(T,0)$. This paper applies high-order multistep methods to discretize an LPTV system eq.˜1, for efficient and accurate computation of its Floquet multipliers and corresponding subspace.
@@
The LPTV systems find applications in a wide range of areas, including dynamic simulation [1], communications [2], control theory [3, 4] and Radio Frequency circuit simulation [5]. A central problem in these fields is the stability analysis of limit cycles in nonlinear autonomous systems. The stability is determined by the variational equation along the cycle. Since the limit cycle is $T$-periodic, its associated Jacobian $G(t)$ is also $T$-periodic, making the variational equation an LPTV system. Floquet theory then provides the framework: the stability is completely characterized by the Floquet multipliers and eigenvectors of this system [6].
Depending on different application background and goals, there are two common scenarios in practice. First, when the matrix $G(t;\delta)$ is given as an explicit function of time $t$, with $\delta$ as a parameter, tracking how Floquet multipliers evolve with respect to this parameter provides insight into the system’s stability. In such cases, a multiplier crossing the unit circle from inside to outside typically signals the onset of a bifurcation or instability [6]. Second, when $G(t)$ is only available at discrete time samples, without an explicit analytic form, the analysis often focuses on computing the Floquet eigenvector associated with the dominant multiplier. In applications such as circuit oscillator design, this largest multiplier is typically equal to one [5]. The corresponding left Floquet eigenvector, known as the perturbation projection vector (PPV) [5, 7], characterizes the system’s phase sensitivity to external perturbations and plays a key role in modeling phase noise in Radio Frequency oscillator systems [5, 8].
@@
The computation of Floquet multipliers and eigenvectors generally involves two key steps: discretization and solving the resulting eigenvalue problem. Existing high-precision software packages such as AUTO-07p [9] and MATCONT [10] adopt high-order one-step collocation methods for discretization, resulting in a periodic eigenvalue problem (periodic-EVP) [6], which can be solved by periodic QR algorithm [11, 12]. This framework makes them well-suited for the first scenario, where the parameterized periodic matrix function $G(t;\delta)$ is explicitly defined. However, the collocation method comes with several limitations for the second scenario, where the system matrix is only available at discrete time samples. It requires over-sampling at Gaussian collocation points within each subinterval, explicit construction of large system matrices, and a condensation step to eliminate over-sampled collocation points. This workflow is computationally expensive and memory-intensive for large-scale systems, and it is not well-suited for efficiently computing Floquet eigenvectors [6].
Motivated by these limitations, and aiming to address both scenarios, this paper applies a multistep method to discretize eq.˜1 that induces a polynomial periodic eigenvalue problem (periodic-PEP)
$\left(\theta^{d}+\sum_{j=0}^{d}\mathcal{A}_{j}\mathcal{C}^{d-j}\theta^{j}\right)u=0,$
where $\mathcal{A}_{j}\mathcal{C}^{d-j}$ are block cyclic matrices constructed from the discretization method. Compared to traditional collocation-based methods, our approach does not require an explicit expression of the system, avoids sampling at fixed Gaussian collocation points, eliminates the need for condensation, and does not explicitly construct or store large coefficient matrices. It naturally supports the simultaneous computation of Floquet multipliers and their eigenvectors, making it especially suitable for large-scale and data-driven systems.
The main contributions of this paper are as follows:
@@
    We apply $d$-step method to an LPTV system, corresponding to a periodic-PEP of degree $d$. We find that the $d$-step discretization generates d1d-1 spurious LPTV systems, whose designed $d$-step discretization corresponds to the same periodic-PEP.
  2. 2.
    We establish convergence analysis based on subspace perturbation theory, as the maximum stepsize $\Delta t$ goes to 0. Among $nd$ eigenpairs of the periodic-PEP, $n$ of them converge to Floquet eigenpairs at rate $\mathcal{O}(\Delta t^{s})$, where $s$ is the consistency order of the multistep method. The remaining n(d1)n(d-1) eigenvalues are related to spurious LPTV systems and decay to 0.
  3. 3.
    We propose an iterative algorithm pTOAR to solve the dominant eigenpairs of the periodic-PEP, and analyze its computational and memory costs. In particular, we show that using high-order multistep method incurs only negligible additional cost compared to the first-order one-step backward Euler method.
The remainder of this paper is organized as follows. In section˜2, the necessary background on Floquet theory, one-step collocation methods, multistep methods, and subspace perturbation theory are reviewed. In section˜3, we analyze how applying a multistep discretization to an LPTV system gives rise to a periodic-PEP. In section˜4, we introduce the spurious systems used to interpret the eigenpairs of the periodic-PEP and define the target objects to which they are expected to converge. In section˜5, the convergence results are established for the eigenpairs using tools from subspace perturbation theory. In section˜6, we present the proposed pTOAR algorithm for computing the dominant eigenpairs of large-scale periodic-PEP, together with an analysis of its computational and memory costs. Numerical experiments are reported in Section˜7, and conclusions are drawn in Section˜8.
2 Preliminaries
2.1 Floquet Theory via the State Transition Matrix of LPTV Systems
Our goal is to compute Floquet multipliers and eigenvectors, defined as the eigenpairs of the state transition matrix (STM) over one period. To this end, we first recall the definitions of these quantities and their key properties, which constitute the core of Floquet theory. The comprehensive Floquet theory can be found in [13, Chapter 5] and [14, Chapter 2].
A typical LPTV system with period $T$ is defined as eq.˜1. Let $X(t)$ denote a fundamental solution matrix, defined by $\dot{X}(t)=G(t)X(t)$, where the initial value matrix $X(0)=X_{0}$ is any full-rank matrix. The associated STM from $s$ to $t$ is $\Phi(t,s)=X(t)X(s)^{-1}$. Note that the definition of $\Phi(t,s)$ is independent of the particular choice of $X_{0}$.
The periodicity of the system matrix $G(t)$ implies that the STM $\Phi(t+T,t)$ is similar to $\Phi(T,0)$ for any $t$, as expressed in eq.˜2:
$\Phi(t+T,t)=\Phi(t,0)\,\Phi(T,0)\,\Phi^{-1}(t,0).$ (2)
To establish eq.˜2, we first note the periodic shift property of the STM,
$\Phi(t+T,T)=\Phi(t,0),\qquad\forall t\in\mathbb{R},$ (3)
which can be proven as follows. Define $Y(t)\equiv X(t+T)\,X^{-1}(T)\,X_{0},$ where $X(t)$ is the fundamental solution matrix. Then $\dot{Y}(t)=G(t)Y(t)$ and $Y(0)=X_{0}$. By the uniqueness of solutions to linear ODEs, we have $Y(t)\equiv X(t)$, which proves eq.˜3. Next, by decomposing $\Phi(t+T,0)$ at $T$ and $t$, we have
$\Phi(t+T,t)\Phi(t,0)=\Phi(t+T,T)\Phi(T,0).$
Right multiplying $\Phi(t,0)^{-1}$ on both sides and then using eq.˜3 yields eq.˜2.
The similarity relation eq.˜2 implies that all matrices $\Phi(t+T,t)$ share the same eigenvalue matrix $R$, whose diagonal entries are the Floquet multipliers. Moreover, there exists a $T$-periodic matrix $U(t)$ such that
$\Phi(t+T,t)=U(t)RU^{-1}(t),$ (4)
where the columns of $U(t)$ and the rows of $U^{-\mathrm{H}}(t)$ provide the corresponding right and left Floquet eigenvectors, respectively.
To construct $U(t)$ explicitly, we should use the the eigen-decomposition at $t=0$, that $\Phi(T,0)=U(0)RU^{-1}(0),$ and write $R=\exp(MT)$, where the diagonal entries of $M$ are the Floquet exponents. Now, we define $U(t)$ as:
$U(t)=\Phi(t,0)U(0)\exp(-Mt).$ (5)
In summary, Floquet multipliers and eigenvectors are obtained from the eigenvalue problem
$\Phi(T,0)u(0)=u(0)\rho,\quad\rho=e^{\mu T},$ (6)
at $t=0$, while the Floquet eigenvectors at any other time $t$ can be recovered using eq.˜5. Therefore, a single eigen-decomposition of $\Phi(T,0)$ suffices to determine all Floquet quantities.
2.2 One-step Methods for Solving Floquet Eigenpairs
Now, to numerically solve eq.˜6, we recall the existing one-step discretization of eq.˜1. The period $[0,T]$ is partitioned at time points $0=t_{0}<t_{1}<\cdots<t_{p}=T$, with stepsizes $\Delta t_{i}=t_{i}-t_{i-1}$. In general, we allow the stepsizes to be variable, and the maximum stepsize is $\Delta t$. Then, we write eq.˜6 on the time grid:
When $\Delta t$ is small enough, one-step methods approximate each $\Phi(t_{i},t_{i-1})$ by a matrix $A^{(i)}$ for $i=1,\ldots,p$, with $u_{i}$ denote the numerical solution at time $t_{i}$. This leads to a periodic eigenvalue problem to compute Floquet eigenpairs:
$A^{(p)}\cdots A^{(2)}A^{(1)}u_{0}=u_{0}\lambda.$ (7)
A well-known one-step method is backward Euler (BE), in which each transition matrix is given by $A^{(i)}=(I-G(t_{i})\Delta t_{i})^{-1}$ when applied to eq.˜1. This method only requires sampling of $G(t_{i})$ at time $t_{i}$. However, its first-order accuracy is usually insufficient for precise Floquet analysis.
High-order one-step methods, such as collocation, improve discretization accuracy. Existing public software packages for computing and analyzing limit cycles, such as MATCONT [10] and AUTO-07p [9], use collocation to discretize periodic systems, and then form the discrete local state transition matrices $A^{(i)}$ to compute Floquet multipliers. In what follows, we briefly outline this procedure; a comprehensive process can be found in [6].
Consider an interval $t_{i-1},t_{i}]$, we first choose the number of collocation points $k$, where a larger $k$ corresponds higher local approximation accuracy. Collocation methods introduce internal stage variables at $k$ equidistant nodes and constructs a Lagrange polynomial interpolant through these points to approximate the solution $x(t)$. The interpolant is then required to satisfy eq.˜1 at $k$ Gaussian collocation points within $t_{i-1},t_{i}]$, which needs sampling the system matrix $G(t)$ at these addtional Guassian collocation points, leading to oversampling.
The next step, known as condensation, is eliminating the intermediate stage variables to obtain the one-step relation $x_{i}=A^{(i)}x_{i-1}$, typically via an LU elimination [6]. This process yields an explicit construction of the local STM $A^{(i)}$.
Although collocation methods have high numerical accuracy, their limitations become apparent depending on the application scenario. In the first scenario, where $G(t)$ is analytically known, and its dimension is small, the overhead of collocation is manageable. However, in the second scenario, where $G(t)$ is only available on given time grids, the collocation methods become both impractical and costly:
  1. 1.
@@ 2.2 One-step Methods for Solving Floquet Eigenpairs
    they rely on temporal internal sampling on Gaussian points of each interval for accuracy, which is expensive and often infeasible in large-scale applications;
@@
2.3 Multistep Methods
We now introduce multistep methods, a class of discretization schemes that attain high accuracy by reusing previously sampled values at adjacent grid points. We recall their general form and the standard definitions of consistency and stability for multistep methods [15, 16].
Definition 1 (Multistep Methods [15][Chapter 3.5] ).
For a fixed time interval $[0,T]$, we still consider a time grid $0=t_{0}<\cdots<t_{p}=T$. For an ODE $\dot{x}=f(t,x)$, a $d$-step method updates the solution at $t_{i}$ by combining the past $d$ steps as:
$\sum_{j=0}^{d}\alpha_{j,i}x_{i-d+j}=\Delta t_{i}\left(\sum_{j=0}^{d}\beta_{j,i}f(t_{i-d+j},x_{i-d+j})\right),$ (8)
where the coefficients $\alpha_{j,i}$ and $\beta_{j,i}$ depend on the past d1d-1 local stepsize ratios $\omega_{i},\omega_{i-1},\ldots,\omega_{i-d+1}$, with $\omega_{i}=\Delta t_{i}/\Delta t_{i-1}$. We normalize $\alpha_{d,i}=1$ and require $\left\lvert\alpha_{0,i}\right\rvert+\left\lvert\beta_{0,i}\right\rvert\neq 0$ to ensure all $d$ steps contribute.
For detailed derivations of these coefficients, see [15, Chapter 3]. As an illustration, the second-order BDF method (Gear2) has coefficients given by eq.˜9.
$\alpha_{0,i}=\frac{{\omega_{i}}^{2}}{1+2\omega_{i}},\,\alpha_{1,i}=-1-\alpha_{0,i},\alpha_{2,i}=1,\,\beta_{2,i}=\frac{1+\omega_{i}}{1+2\omega_{i}},\,\beta_{1,i}=\beta_{0,i}=0.$ (9)

### v2

@@ Abstract
Accurate and efficient computation of Floquet multipliers and subspaces is essential for analyzing limit cycle in dynamical systems and periodic steady state in Radio Frequency simulation. This problem is typically addressed by solving a periodic linear eigenvalue problem, which is discretized from the linear time-periodic system using one-step collocation methods. Collocation methods become costly for large-scale cases. Our alternative approach is to use multistep methods. The multistep method leads to a periodic polynomial eigenvalue problem (pPEP), and introduces additional parasitic periodic eigenvalues. We prove that as the stepsize decreases, the computed Floquet multipliers and their associated invariant subspace converge with higher order, while the parasitic periodic eigenvalues converge to zero geometrically and therefore Floquet multipliers are not affected by those parasitic ones. A memory-efficient algorithm pTOAR is designed to solve the large-scale pPEP. Its computational and memory costs are almost independent of the choice of multistep methods. Numerical results coincide with our convergence analysis, and also demonstrate the efficiency of pTOAR.
@@ 1 Introduction
A linear time-periodic (LTP) system is described by
(1) $\frac{\mathrm{d}x(t)}{\mathrm{d}t}=G(t)x(t),$
@@
where $G(t)\in\mathbb{R}^{n\times n}$ is a $T$-periodic matrix function, i.e., $G(t+T)=G(t)$, and $x(t)\in\mathbb{R}^{n}$ is the state vector. The monodromy matrix of Eq.˜1 is obtained by integrating Eq.˜1 from 0 to $T$ with initial value matrix $X(0)=I_{n}$. The Floquet multipliers and eigenvectors are the eigenpairs of the monodromy matrix. Efficient and accurate computation of Floquet multipliers and corresponding eigenvectors is essential for analyzing LTP systems.
@@
The LTP systems find applications in a wide range of areas, including dynamic simulation [1], communications [15], control theory [6, 4] and Radio Frequency circuit simulation [9]. The central problem in these fields is the stability analysis for the limit cycles of nonlinear autonomous systems. The stability is completely characterized by the Floquet multipliers and eigenvectors of the variational equation along the limit cycle in the form of Eq.˜1 [20].
@@
The computation of Floquet multipliers and eigenvectors generally involves two key steps: discretization of Eq.˜1 and solving the resulting eigenvalue problem. The simplest implementation is to use backward Euler (BE) method to obtain a periodic eigenvalue problem (pEVP), which can be solved by periodic QR algorithm [5, 18]. Existing high-precision software packages such as AUTO-07p [12] and MATCONT [11] adopt high-order one-step collocation methods for discretization. Collocation methods require extra function evaluations at Gaussian collocation points within each subinterval, and a condensation step to eliminate the extra collocation points. This framework makes collocation well-suited for modest-scale cases whose coefficient function $G(t)$ is given by an analytical expression. For large-scale systems, collocation methods become computationally expensive and memory-intensive [20], particularly as the condensation may break the inherent sparsity of the systems. In some practical cases, collocation methods also struggle if extra function evaluations are not permitted.
This paper applies multistep methods to discretize Eq.˜1 that induces a periodic polynomial eigenvalue problem (pPEP). Multistep methods improve accuracy under the same sampling used by BE. However, applying a $d$-step method, the resulting pPEP possesses $nd$ periodic eigenpairs under cyclic symmetry; and the pPEP is much more complicated to solve.
By introducing parasitic LTP systems, we prove that n(d1)n(d-1) periodic eigenvalues decay to 0 at rate $\mathcal{O}(\Delta t)$ as the maximum stepsize $\Delta t$ goes to 0. The remaining $n$ eigenpairs converge to Floquet eigenpairs at rate $\mathcal{O}(\Delta t^{s})$, where $s$ is the consistency order of the multistep method. Therefore, we only need to solve the $n$ dominant periodic eigenvalues of pPEP.
To efficiently solve the eigenpairs of the pPEP, we propose the pTOAR. In particular, we show that using high-order multistep method incurs only negligible additional cost compared to the first-order one-step backward Euler method.
The remainder of this paper is organized as follows. In Section˜2, the necessary background on Floquet theory, one-step collocation methods, multistep methods, and subspace perturbation theory are reviewed. In Section˜3, we discuss the formulation of pPEP. In Section˜4, the convergence results are established for the eigenpairs using tools from subspace perturbation theory. In Section˜5, we present the pTOAR for solving pPEP, together with an analysis of its computational and memory costs. Numerical experiments are reported in Section˜6, and conclusions are drawn in Section˜7.
2 Preliminaries
2.1 Floquet Theory
Floquet theory provides the structure of the solutions of LTP systems Eq.˜1. We recall the key results of Floquet theory, and the comprehensive version can be found in [2, Chapter 5] and [14, Chapter 2].
For LTP system Eq.˜1 with period $T$, let $X(t)$ denote a fundamental solution matrix, defined by $\dot{X}(t)=G(t)X(t)$, where the initial value matrix $X(0)=X_{0}$ is any nonsingular matrix. Define the state transition matrix $\Phi(t,s)$ by as the mapping from the principal fundamental solution matrix at time $s$ to that at time $t$, that $\Phi(t,s)X(s)=X(t)$.
Theorem 2.1 (Floquet Representation Theorem [2][Theorem 20.9]).
For LTP system Eq.˜1, there exists a nonsingular constant matrix $M$ and a $T$-periodic matrix function $U(t)\in C^{1}(\mathbb{R},\mathbb{C}^{n^{2}})$ such that its state transition matrix $\Phi(t,0)$ can be written as
(2) $\Phi(t,0)=U(t)e^{Mt}.$
Thus, the state transition matrix over a period $\Phi(t+T,t)$ can be decomposed by:
(3) $\Phi(t+T,t)=U(t)e^{MT}U^{-1}(t),$
Since $M$ is a constant matrix, and $T$ is the period, all matrices $\Phi(t+T,t)$ are similar. The eigenvalues of $M$ and $e^{MT}$ are called Floquet exponent and Floquet multipliers. For simplicity, we assume all Floquet multipliers are semi-simple. Thus, we assume $e^{MT}$ is diagonal, and Eq.˜3 becomes an eigenvalue decomposition of $\Phi(t+T,t)$. The columns of $U(t)$ and the rows of $U^{-\mathrm{H}}(t)$ provide the corresponding right and left Floquet eigenfunctions, respectively.
To obtain the Floquet multipliers, we only need to consider the eigenvalue problem Eq.˜4, which corresponds to Eq.˜3 at $t=0$:
(4) $\Phi(T,0)u(0)=u(0)e^{\mu T}.$
$u(0)$ is called the Floquet eigenvector associated with the Floquet multiplier $\lambda=e^{\mu T}$.
Though Eq.˜4 is only a slice of Eq.˜3 at $t=0$, it is sufficient to solve Floquet eigenfunctions $U(t)$ within the period. Once all the Floquet multipliers $e^{MT}$ are obtained, we can figure out all the Floquet exponents $M$, and thus solve Floquet eigenfunctions $U(t)$ from Eq.˜2. We should also note that $u(0)$ in Eq.˜4 is the value of the Floquet eigenfunction at $t=0$.
2.2 One-step Methods for Solving Floquet Eigenpairs
Let the period $[0,T]$ be partitioned at time points $0=t_{0}<t_{1}<\cdots<t_{p}=T$, with stepsizes $h_{i}=t_{i}-t_{i-1}$ allowed to be variable, and with maximum $h$. Rewrite Eq.˜3 on the time grid:
When $h$ is small enough, one-step methods approximate each $\Phi(t_{i},t_{i-1})$ by a matrix $A^{(i)}$ for $i=1,\ldots,p$, and approximate the eigenpair $\left(u(t_{0}),e^{\mu T}\right)$ by $\left(u_{0},\lambda\right)$. This leads to a periodic eigenvalue problem [18]:
(5) $A^{(p)}\cdots A^{(2)}A^{(1)}u_{0}=u_{0}\lambda.$
One of the simplest one-step method is backward Euler (BE). BE only requires function evaluation of $G(t_{i})$ at time $t_{i}$, and $A^{(i)}=\left[I-G(t_{i})h_{i}\right]^{-1}$ when applied to Eq.˜1. However, its first-order accuracy is usually insufficient in practical.
High-order one-step methods, such as collocation, improve discretization accuracy. In what follows, we briefly outline this procedure; a comprehensive process can be found in [20].
The first step is to choose the number of collocation points $k$ within an interval $t_{i-1},t_{i}]$, where a larger $k$ corresponds higher local approximation accuracy. Collocation methods introduce internal stage variables at $k$ equidistant nodes and constructs a Lagrange polynomial interpolant through these points to approximate the solution $x(t)$. The interpolant is then required to satisfy Eq.˜1 at $k$ Gaussian collocation points within $t_{i-1},t_{i}]$, which needs extra function evaluation of the system matrix $G(t)$ at these additional Gaussian collocation points.
The next step, known as condensation, is eliminating the intermediate stage variables to obtain the one-step relation $x_{i}=A^{(i)}x_{i-1}$, typically via an LU elimination [20]. This process yields an explicit construction of the local state transition matrix $A^{(i)}$.
Collocation methods achieve higher accuracy than BE, and are efficient for cases that $G(t)$ is analytically known. It is also adopted by existing mature public software packages for computing and analyzing limit cycles, such as MATCONT [11] and AUTO-07p [12]. However, if $G(t)$ is only available on given time grids, the collocation methods become both impractical and costly:
@@ 2.2 One-step Methods for Solving Floquet Eigenpairs
    they rely on temporal internal function evaluation on Gaussian points of each interval for accuracy, which is expensive and often infeasible in large-scale applications;
@@
2.3 Multistep Methods for Solving Initial Value Problems
Multistep methods offer an alternative approach to enhance discretization accuracy by utilizing previously computed values at neighboring grid points. We review their general form and recall standard notions of consistency and stability [17]. These results are essential for solving initial value problems, and also provide the foundation for our later analysis of Floquet multipliers using multistep discretizations.
Definition 2.2 (Variable Stepsize Multistep Methods [17, Chapter 3.5]).
Let $0=t_{0}<\cdots<t_{p}=T$ be a time grid on the interval $[0,T]$. We define the stepsize $h_{i}=t_{i}-t_{i-1}$ and the stepsize ratio $\omega_{i}=h_{i}/h_{i-1}$. We define the ratio vector $\boldsymbol{\omega}_{i}\in\mathbb{R}^{d}$ as: $\boldsymbol{\omega}_{i}=(\omega_{i-d+1},\ldots,\omega_{i}).$ For an ODE $\dot{x}=f(t,x)$, a $d$-step method updates the solution at $t_{i}$ according to:
(6) $\sum_{j=0}^{d}\alpha_{j}^{(i)}x_{i-d+j}=h_{i}\sum_{j=0}^{d}\beta_{j}^{(i)}f(t_{i-d+j},x_{i-d+j}),$
where the coefficients are determined by smooth functions of the local step ratios, i.e., $\alpha_{j,i}=\alpha_{j}(\boldsymbol{\omega}_{i})$ and $\beta_{j,i}=\beta_{j}(\boldsymbol{\omega}_{i})$, and $\left\lvert\alpha_{0,i}\right\rvert+\left\lvert\beta_{0,i}\right\rvert\neq 0$ for irreducibility condition.
For detailed derivations of these coefficients, see [17, Chapter 3]. As an illustration, the second-order BDF method (Gear2) has coefficients given by Eq.˜7.
(7) $\alpha_{0,i}=\frac{{\omega_{i}}^{2}}{1+2\omega_{i}},\,\alpha_{1,i}=-1-\alpha_{0,i},\alpha_{2,i}=1,\,\beta_{2,i}=\frac{1+\omega_{i}}{1+2\omega_{i}},\,\beta_{1,i}=\beta_{0,i}=0.$